Operations Report – January 22nd

Crew 187 Operations Report 22012018

SOL: 10

Name of person filing report: Luis Jose Diaz Lopez

Non-nominal systems: None.

Notes on non-nominal systems: Generator system limping along with a now-nominal routine.

Generator (hours run): 15h 00min

Generator turned off, charging battery at 10h:00m

Generator turned on at 19h:30m

Solar— SOC

@ 08:38: 53%

@ 19:35: 43%

Diesel: It can’t be measure. Explained in the Report 19JAN2018

Propane: 92%

Ethanol Free Gasoline (5 Gallon containers for ATV): 4.75 Gallons

Water (trailer): 0 Gallons

Water (static): 100 Gallons

Trailer to Static Pump used: Yes

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 01301216 Gallons

Toilet tank emptied: No

ATVs Used: 350(1), 350(2)

Oil Added? No

ATV Fuel Used: As a result of today’s EVA incident, I wasn’t able to check the final level of the fuel. Additionally, I don’t know exactly how to do it properly, but I can make an estimation. Tomorrow, before the EVA they will be refilled.

# Hours the ATVs were used today: 2h 40 min

Notes on ATVs: ATVs were nominal.

Deimos rover used: Yes

Hours: 102.3 h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge: –

Currently charging: Maybe

Spirit rover used: Yes

Hours: 12.7h

Beginning charge: 100%

Ending charge: –

Currently charging: No

Opportunity rover used: Yes

Hours: 5.6h

Beginning charge: 100%

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 39h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Water heater was fixed; however, you have to open the pipe handle completely in order to the water heater to work. Kinda stressful because our critical resource is water. It will be great if this issue can be fix. The lower deck heater is now operating.

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: Today a water filter was installed in the water loft.

Summary of EVA operations: During the EVA of today, the battery level of Deimos rover dropped suddenly from 71% to 26% in just 5 mins. Due to this, we had to tow the rover with a ATV until the Hab.

Summary of GreenHab operations: All nominal

Summary of Science Dome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Today the water heater was fixed but a lot of water was wasted to do that (testing and fixing), that’s the reason that today we have just 100 Gallons of water; please, we will need a refill of this. During the towing process the jumping cables were broke, We need a new pair of this.

Copyright © The Mars Society. All rights reserved. | Main Site